Calculus II, Spring 2016 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 02

1. Section 10.1: Problems 53, 59, 63, 67, 69, 81, 87, 97, 98.

2. Section 10.2: Problems 61, 65, 71, 75.

3. Section 10.4: Problems 17, 27, 29, 31, 43, 45, 51, 61, 62. Note in problem 62: the case p = 1, q < −1 should be convergent, instead.

4. Section 10.5: Odd problems in problem 17-43, 61, 65.